# Topological Corner States in Bilayer and Trilayer Systems with Vertically Stacked Topologically Distinct Layers

### v1

@@ Abstract
We investigate bilayer and trilayer systems composed of topologically distinct, vertically stacked layers based on the Benalcazar-Bernevig-Hughes model. We have identified a topological phase transition that significantly alters the number of the topological corner states in these systems. Additionally, we find that traditional nested Wilson loop analysis inaccurately classifies certain phases, leading us to evaluate multipole chiral numbers (MCNs) as a more appropriate topological invariant. This approach accurately identifies topological phases and the number of MCNs coincides with the number of corner modes, enhancing our understanding of topological insulators and opening avenues for further applications.
@@ I Introduction
Topological insulators are fascinating emerging materials found in condensed matter physics[1, 2, 3]. These materials are distinguished by their unique combination of an insulating bulk and metallic surfaces, i.e., edge states. The distinctive behavior of the edge states originates from the topological features in the bulk band structure giving a non-zero topological invariant, demonstrating the essential concept of bulk-boundary correspondence[4, 5]. The basic concept of topological materials has been expanded into classical wave phenomena, including photonics[6, 7, 8, 9, 10], acoustic waves[11, 12, 13, 14], mechanical waves[15, 16, 17, 18], thermal waves[19], as well as electric circuits[20, 21, 22, 23, 24]. This expansion has led to attractive applications such as topological lasers[25, 26].
@@
Expanding beyond the conventional topological insulators, the concept of a higher-order topological insulators (HOTIs [27, 28]) has been proposed and is gaining significant attention. This class of materials is characterized by topological states that are at least two dimensions lower than the system itself. For instance, in a two-dimensional (2D) HOTI, one can observe zero-dimensional corner states.
@@
Building on this framework, the majority of research has been directed towards the study of 2D monolayers[29, 30, 31, 32] and their vertically stacked three-dimensional structures[33, 34, 35, 36, 37, 38]. These investigations have a broad scope, aiming to reveal the distinct topological properties intrinsic to these structures[39, 40, 41], as well as to examine their potential applications in diverse fields, including nanocavity lasers [42, 43, 44], nonlinear optics[45, 46], cavity quantum electrodynamics (QED) experiments[47, 48] and interferometer[49, 50]. There are some studies on vertically stacking of identical models in multilayer topological systems [51, 52, 53]. However, multilayer systems composed of topologically distinct layers have not been investigated.
@@
In this paper, we investigate multilayer systems composed of topologically distinct layers. Our work is based on the Benalcazar-Bernevig-Hughes (BBH) model [27], a simplest model for studying quadrupole topological insulators. This model consists of $\pi$-flux square lattices and it supports topological corner modes. The realization of quadrupole insulators has been demonstrated through various experimental setups, including optical ring resonator arrays [54], waveguide arrays[55], electric circuits[21], acoustic system[56], and microwave resonator[57].
@@
Our findings indicate a phase transition induced by the interlayer coupling, which significantly influences the emergence and annihilation of topological corner states. Interestingly, we identified a phase where the conventional method, using the nested Wilson loop, incorrectly classifies the phase as trivial, despite the presence of corner states. Therefore, we calculated a recently developed alternative topological invariant known as multipole chiral numbers (MCNs) [58] to evaluate the topological phases. Due to its capability to handle $\mathbb{Z}$-class higher-order topological systems, the MCN has begun to be applied across various models[59, 60, 61, 62]. We found that the MCN precisely identifies the phases present in our model and accurately matches the number of corner modes at zero-energy. By adopting the MCN as a topological invariant, we not only identify changes in the number of corner modes but also uncover the phase transition induced by interlayer coupling.
@@ II Topological corner states
The BBH lattice consists of four sites in a unit cell where the $\pi$-flux is inserted by introducing the negative coupling in a plaquette, as shown with the red bond in Fig. 1 (a). There are two topologically distinct phases, determined by the ratio of the intra-cell coupling to the inter-cell coupling. When the intra-cell coupling is weaker than inter-cell coupling, there exists four degenerate states that form a single set of corner modes, localized at each corner of the structure within a band gap at energy $E=0$. The bulk Hamiltonian for the system shown in Fig. 1 (a) is given by
@@
Throughout this paper, we use $\kappa_{1}$ ($\kappa_{2}$) and $\kappa_{2}$ ($\kappa_{1}$) as the intra-cell and inter-cell couplings for the topological (trivial) system, respectively, with its corresponding Hamiltonian $\mathcal{H}_{\text{BBH}}^{\text{Topo}}\left(k_{x},k_{y}\right)$ ($\mathcal{H}_{\text{BBH}}^{\text{Tri}}\left(k_{x},k_{y}\right)$). Here, we suppose $\kappa_{1}<\kappa_{2}$, and both are positive.
The BBH system has two occupied degenerate bands that possess two cancelling dipole moments and a non-vanishing quadrupole moment. By adapting the Wilson loop calculation, the degeneracy is lifted, resulting in two distinct single-band subspaces where Wannier bands become non-degenerate. Consequently, each Wannier band carries its own topological invariant, which can be obtained through the nested Wilson loop calculation [27]. Following the calculation, a quadrupole moment $q_{xy}=\frac{1}{2}$ signifies that the system is topological, while $q_{xy}=0$ signifies that the system is trivial.
In our study, we consider a vertically stacked configuration composed of topologically distinct layers, taking into account the vertical nearest-neighbor site coupling denoted by $t_{\mathrm{v}}$. Fig. 1 (b) illustrates the considered BBH bilayer system and its unit cell, where the top and bottom layers are topological and trivial, respectively. For trilayer systems, we consider two cases as shown in Figs. 1 (d) and (e), where the arrangements of the topological and trivial layers are flipped.
II.1 Topological corner states in BBH bilayer lattice
The Hamiltonian describing the bilayer system, shown in Fig. 1 (b), is given by
@@ II.1 Topological corner states in BBH bilayer lattice
Figure 1: (a) Blue and red bonds represent the positive and negative couplings, respectively. Corner states, shown in red at each corner, appear when $\kappa_{1}<\kappa_{2}$. (b) An illustration of the BBH bilayer system composed of topological and trivial layers. $\kappa_{1}$ ($\kappa_{2}$) and $\kappa_{2}$ ($\kappa_{1}$) denote the intra- and inter-cell couplings for topological (trivial) system, respectively.
(c) The labeling lattice sites a, b, c, and d represent the Hamiltonian basis for each layer. (d,e) Two distinct BBH trilayer systems where topologically distinct layers are stacked on top of each other.
@@
Here, we use the site index shown in Fig. 1(c) in which chiral symmetry is preserved. This ordering is essential for the theoretical analysis presented later in this paper. $\sigma_{x}$ and $\mathbb{I}_{2}$ are Pauli matrix and two by two identical matrix, respectively. The band structure has four pairs of doubly degenerate bands, and exhibits both opening and closing of a band gap around the $\Gamma$ point as $t_{\mathrm{v}}$ changes. (Fig. 2). The critical value $t_{\mathrm{c}}^{\rm{bilayer}}$ corresponds to the point at which the band gap closes, defined as $t_{\mathrm{c}}^{\rm{bilayer}}=\sqrt{2}(\kappa_{1}+\kappa_{2})$. See appendix A for details.
@@
Figure 3: (a) Energy spectrum vs $t_{\mathrm{v}}$ for BBH bilayer system. Blue dots represent the bulk spectrum, with or without gapped edges, whereas the red line corresponds to a single set of corner modes. The gapped edge states appear when corner modes exist. (b) Field amplitude summed over the four corner modes at $E=0$ for $t_{\mathrm{v}}$=$t_{\mathrm{c}}^{\rm{bilayer}}-1$. Parameters are chosen as $\kappa_{1}$=1, $\kappa_{2}=3$ and $N=20$.
We now consider a finite BBH model with square geometry, with each layer forming a BBH lattice consisting of $2N\times 2N$ sites. Fig. 3(a) shows the energy spectrum as a function of varying $t_{\mathrm{v}}$. In the weak interlayer coupling regime ($t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{bilayer}}$), the system possesses a single set of corner modes, consisting of a total of four degenerate states, each predominantly localized at the corners of the layers. These are depicted by the red line in the figure. Fig. 3(b) shows the field amplitude of the corner state, summed over for all four states. Notably, these corner states disappear when the interlayer coupling exceeds the critical value $t_{\mathrm{c}}^{\rm{bilayer}}$. Sudden disappearance of the corner states, accompanied by a gap closing, suggests a topological phase transition associated with increasing interlayer coupling strengths.
In order to determine the topological phases, we calculate the quadrupole moment $q_{xy}$ [27] via the nested Wilson loop. As a result, we find that $q_{xy}=\frac{1}{2}$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{bilayer}}$, identifying the system as topological, while $q_{xy}=0$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{bilayer}}$, indicating a trivial phase. These results indicate that the interlayer coupling induces a topological phase transition in the BBH bilayer system. The presence and absence of topological corner states change at the critical value $t_{\mathrm{c}}^{\rm{bilayer}}$ bilayer as the band gap closes and re-opens. This critical value marks the transition point where the number of topological corner states changes. The detailed calculation of the quadrupole moment is shown in the Appendix B.
II.2 Topological corner states in BBH trilayer lattice
We proceed to study a BBH trilayer system depicted in Fig. 1(d). This arrangement consists of a trivial layer positioned between two topological layers, where the layers are coupled by an interlayer coupling $t_{\mathrm{v}}$.
@@ II.2 Topological corner states in BBH trilayer lattice
with the interlayer coupling matrix $\mathcal{H}_{\text{v}}$. Similar to the bilayer system, the labeling of sites in the Hamiltonian varies between odd and even layers to maintain the chiral symmetry, as shown in Fig. 1(c). This labeling is crucial for applying the methods we will discuss in the next section. The trilayer system possesses six pairs of doubly degenerate bands. Analogous to the bilayer system, band gap closes at a critical value $t_{\mathrm{c}}^{\rm{trilayer}}=\kappa_{1}+\kappa_{2}$ (See appendix A). The band gap re-opens again when $t_{\mathrm{v}}$ exceeds the critical value.
@@
Figure 4: (a) Energy spectrum vs $t_{\mathrm{v}}$ for BBH trilayer system (topological-trivial-topological structure). Blue dots represent the gapped edge states and bulk spectrum. The green and red lines correspond to four and a single set of corner modes, respectively. (b) Field amplitude summed over the four corner modes at $E=0$ for $t_{\mathrm{v}}$=$t_{\mathrm{c}}^{\rm{trilayer}}-1$ and $t_{\mathrm{v}}$=$t_{\mathrm{c}}^{\rm{trilayer}}+1$. ’Corner 1’ persists even after exceeding the critical value, while ’corner 2’ vanishes at the critical value $t_{\mathrm{c}}^{\rm{trilayer}}$. Parameters are chosen as $\kappa_{1}$=1, $\kappa_{2}=3$, and $N=20$.
@@
For a finite BBH trilayer system with square geometry, with each layer consisting of $2N\times 2N$ sites, we calculate the eigenenergies as a function of the interlayer coupling strength(Fig 4(a)). When $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$, which is below the critical value, there are two sets of corner states, with a total number of eight (illustrated by the green line in the figure). In contrast, when $t_{\mathrm{v}}$ exceeds $t_{\mathrm{c}}^{\rm{trilayer}}$, only a single set of corner states survives, with a total number of four (indicated by the red line). Each of these sets consists of four degenerate modes located at $E=0$. These two sets of corner states are distinguished by their unique field distributions(Fig. 4 (b)). Specifically, the corner states that disappear at the critical interlayer coupling show non-zero field amplitudes across all layers, which corresponds to symmetric coupling between the corners (corner 2). Conversely, the corner states that remain after exceeding the critical interlayer coupling exhibit a complete absence of field amplitude in the middle layer, which corresponds to anti-symmetric coupling between the corners (corner 1). The presence of topological corner states, ’Corner 1’, at large interlayer couplings, is understood by an effective model in the vicinity of the zero-energy, derived based on the perturbation theory. See Appendix C.
Figure 5: (a) Energy spectrum vs $t_{\mathrm{v}}$ for BBH trilayer system (trivial-topological-trivial structure). Blue dots correspond to the bulk spectrum, with or without gapped edges, whereas the red line represents a single set of corner modes. Gapped edges are present when the system exhibits corner states. (b) Field amplitude summed over the four corner modes at $E=0$ for $t_{\mathrm{v}}$=$t_{\mathrm{c}}^{\rm{trilayer}}-1$. Parameters are chosen as $\kappa_{1}$=3, $\kappa_{2}=1$ and $N=20$.
@@
As discussed in the subsection $\mathrm{II.A}$ of the bilayer system, we calculate the quadrupole moment $q_{xy}$ to identify the topological phases of the trilayer system. Consequently, our analysis shows that $q_{xy}=0$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$, classifying the system as trivial, despite the presence of two sets of corner states at $E=0$. In the case where $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$, we have $q_{xy}=\frac{1}{2}$, indicating the system as topological. This leads to the issue where the quadrupole moment fails to fully capture the topological phases in certain instances. Therefore, the introduction of another invariant is essential to accurately describe the topological phases of both bilayer and trilayer BBH systems. We will calculate and discuss this alternative invariant in the next section.
@@
Next, we study a trilayer system which involves a topological layer sandwiched between two trivial layers, with each layer coupled by an interlayer coupling $t_{\mathrm{v}}$. This arrangement is depicted in Fig. 1 (e). In this case, the Hamiltonian is given by
@@
Fig. 5(a) shows the energy spectrum as a function of $t_{\mathrm{v}}$. There is a single set of corner modes that includes four degenerate corner states at $E=0$ (depicted by the red line) when $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$; however, all corner modes disappear above the critical value $t_{\mathrm{c}}^{\rm{trilayer}}$. The field amplitudes of the corner modes, shown in Fig. 5 (b), are distributed across each layer. The quadrupole moments are obtained as $q_{xy}=\frac{1}{2}$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$, identifying the phase as topological, while $q_{xy}=0$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$, classifying it as trivial. We note that in BBH trilayer systems with large interlayer coupling $t_{\mathrm{v}}$, the size of the band-gap varies between two distinct systems, despite them having identical band structures, as shown in Figs 4(a) and 5(a). This discrepancy arises due to the presence of gapped edge states within the band gap as illustrated in Fig. 4(a).

### v2

# Topological Corner States in Bilayer and Trilayer Systems with Vertically Stacked Topological Heterostructures

@@ Abstract
We investigate bilayer and trilayer systems composed of topologically distinct, vertically stacked layers, forming topological heterostructures based on the Benalcazar-Bernevig-Hughes model. We find that a topological phase transition induced by interlayer coupling significantly alters the number of corner states in these topological structures. Furthermore, we find that traditional nested Wilson loop analysis inaccurately classifies certain phases, leading us to evaluate multipole chiral numbers (MCNs) as a more appropriate topological invariant for this scenario. The MCNs not only enable accurate classification of topological phases but also directly correspond to the number of zero-energy corner states, effectively characterizing $\mathbb{Z}$-class HOTI phases. Our study proposes the novel concept of topological heterostructures, providing critical insights into the control of localized corner states within multilayer systems and expanding potential research directions.
@@ I Introduction
Topological insulators have emerged as fascinating materials in condensed matter physics[1, 2, 3], characterized by an insulating bulk and conductive edge states. These edge states arise from the topological features in the bulk band structure, which result in a non-zero topological invariant, thus demonstrating the concept of bulk-boundary correspondence[4, 5]. The basic principles of topological materials have been extended beyond condensed matter systems to classical wave phenomena, such as photonics[6, 7, 8, 9, 10], acoustic waves[11, 12, 13, 14], mechanical waves[15, 16, 17], thermal waves[18], and electric circuits[19, 20]. This expansion has led to attractive applications such as topological lasers[21, 22] and topological electronic devices[23, 24].
@@
In recent years, the concept of higher-order topological insulators (HOTIs)[25, 26] has gained significant attention, expanding the scope of topological materials. HOTIs are characterized by topological states that are at least two dimensions lower than that of the system itself; for example, a two-dimensional (2D) HOTI can host zero-dimensional corner states.
@@
Building on this framework, the majority of research has been directed towards the study of 2D HOTIs[27, 28, 29, 30, 31, 32, 33, 34, 35] and their three-dimensional vertically stacked extensions, which can host hinge states[36, 37, 38, 39, 40, 41]. These studies have aimed to uncover the unique topological properties of these systems and explore their potential applications in various fields, such as nanocavity [42, 43], topological nanolasers[44, 45], nonlinear optics[46, 47], and cavity quantum electrodynamics (QED)[48, 49].
For multilayer systems, those composed of vertically stacked layers, known as van der Waals heterostructures [50], have been extensively studied. These systems, characterized by weak interlayer coupling relative to in-plane coupling, have demonstrated unique topological properties[51, 52] and show promise for applications in various fields[53, 54]. Despite this progress, the investigation of multilayer systems consisting of topologically distinct layers remains largely unexplored.
@@
In this paper, we investigate multilayer systems composed of topologically distinct layers, forming topological heterostructures. We focus on the effects of corner states and topological phases in relation to interlayer coupling strength. Our work is based on the Benalcazar-Bernevig-Hughes (BBH) model [25], the simplest model for studying quadrupole topological insulators. This model consists of $\pi$-flux square lattices and supports topological corner states. The realization of quadrupole insulators has been demonstrated through various experimental setups, including optical ring resonator arrays [55], waveguide arrays[56], electric circuits[57], acoustic system[58], and microwave resonator[59].
@@
Our findings indicate a topological phase transition induced by the interlayer coupling, which significantly influences the emergence and annihilation of topological corner states. Interestingly, we identify a phase in which the conventional nested Wilson loop method incorrectly classifies the topological phase as trivial, despite the presence of corner states. Therefore, we calculate an alternative topological invariant, recently developed multipole chiral numbers (MCNs) [60], to evaluate the topological phases. Due to its ability to handle $\mathbb{Z}$-class chiral-symmetric HOTI systems[61, 62], the MCN has been increasingly applied to various models [63, 64, 65, 66, 67, 68]. We found that the MCN precisely identifies the phases present in our model and accurately matches the number of corner states at zero energy. By adopting the MCN as a topological invariant, we not only identify changes in the number of corner states, but also uncover the topological phase transition induced by interlayer coupling.
@@ II Topological corner states
The BBH model consists of four sites in a unit cell where the $\pi$-flux is inserted by introducing the negative coupling in a plaquette, as shown with the red bond in Fig. 1 (a). There are two topologically distinct phases, determined by the ratio of the intracell coupling to the intercell coupling. When the intra-cell coupling is weaker than inter-cell coupling, there exists four degenerate states that form a single set of corner modes, localized at each corner of the structure within a band gap at energy $E=0$. The bulk Hamiltonian for the system shown in Fig. 1 (a) is given by
Figure 1: (a) Blue and red bonds represent the positive and negative couplings, respectively. Corner states, shown in red at each corner, appear when $\kappa_{1}<\kappa_{2}$. (b) the 1st Brillouin zone. (C) Illustration of the BBH bilayer system composed of topological and trivial layers. $\kappa_{1}$ ($\kappa_{2}$) and $\kappa_{2}$ ($\kappa_{1}$) denote the intra- and inter-cell couplings for topological (trivial) system, respectively. (d) The labeling lattice sites a, b, c, and d represent the Hamiltonian basis for each layer. (e,f) Two distinct BBH trilayer systems in which topologically distinct layers are stacked on top of each other.
@@
Throughout this paper, we use $\kappa_{1}$ ($\kappa_{2}$) and $\kappa_{2}$ ($\kappa_{1}$) as intracell and intercell couplings for the topological (trivial) system, respectively, with its corresponding Hamiltonian $\mathcal{H}_{\text{BBH}}^{\text{Topo}}\left(k_{x},k_{y}\right)$ ($\mathcal{H}_{\text{BBH}}^{\text{Tri}}\left(k_{x},k_{y}\right)$). Here, we suppose $\kappa_{1}<\kappa_{2}$, and both are positive.
The BBH system has two occupied degenerate bands that possess two canceling dipole moments and a non-vanishing quadrupole moment. By adopting the Wilson loop calculation, the degeneracy is lifted, resulting in two distinct single-band subspaces where Wannier bands become non-degenerate. Consequently, each Wannier band carries its own topological invariant, which can be obtained through the nested Wilson loop calculation [25]. Following the calculation, a quadrupole moment $q_{xy}=\frac{1}{2}$ indicates that the system is topological, while $q_{xy}=0$ means that the system is trivial.
In this paper, we consider a vertically stacked configuration composed of topologically distinct layers, taking into account the vertical coupling of the nearest neighbor site denoted by $t_{\mathrm{v}}$. Fig. 1 (c) illustrates the considered BBH bilayer system and its unit cell, where the top and bottom layers are topological and trivial, respectively. For trilayer systems, we consider two cases as shown in Figs. 1 (e) and (f), where the arrangements of the topological and trivial layers are flipped.
II.1 Topological corner states in BBH bilayer system
The Hamiltonian describing the bilayer system, shown in Fig. 1 (c), is given by
@@ II.1 Topological corner states in BBH bilayer system
Here, we use the site index shown in Fig. 1(d) in which chiral symmetry is preserved. This ordering is essential for the theoretical analysis presented later in this paper. $\sigma_{x}$ and $\mathbb{I}_{2}$ are the Pauli matrix and two-by-two identity matrix, respectively. The band structure has four pairs of doubly degenerate bands, and exhibits both opening and closing of a band gap around the $\Gamma$ point as $t_{\mathrm{v}}$ changes. (Fig. 2). The critical value $t_{\mathrm{c}}^{\rm{bilayer}}$ corresponds to the point at which the band gap closes, defined as $t_{\mathrm{c}}^{\rm{bilayer}}=\sqrt{2}(\kappa_{1}+\kappa_{2})$. See Appendix A for details.
@@
Figure 3: (a) Energy spectrum vs $t_{\mathrm{v}}$ for BBH bilayer system. Blue dots represent the bulk spectrum, with or without gapped edges, whereas the red line corresponds to a single set of corner modes. The gapped edge states appear when corner modes exist. (b) Field amplitude summed over the four corner modes at $E=0$ for $t_{\mathrm{v}}$=$t_{\mathrm{c}}^{\rm{bilayer}}-1$. The inset shows an enlarged view of a single corner, with the gray area representing a unit cell. Parameters are chosen as $\kappa_{1}$=1, $\kappa_{2}=3$ and $N=20$.
We now consider a finite BBH model with square geometry, with each layer forming a BBH lattice consisting of $2N\times 2N$ sites. Fig. 3(a) shows the energy spectrum as a function of $t_{\mathrm{v}}$. In the weak interlayer coupling regime ($t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{bilayer}}$), the system possesses a single set of corner modes, consisting of a total of four degenerate states, each predominantly localized at the corners of the layers. This is depicted by the red line in the figure. Fig. 3(b) shows the field amplitude of the corner state, summed for all four states. An enlarged view of the corner modes reveals their localization on a single sublattice at each corner, attributed to chiral symmetry. A detailed analysis of the field distribution, considering the limit $\kappa_{1}=0$, is presented in Appendix B. These corner states disappear when the interlayer coupling exceeds the critical value $t_{\mathrm{c}}^{\rm{bilayer}}$. This disappearance can be understood through perturbation theory, as explained in detail in Appendix D. Sudden disappearance of the corner states, accompanied by a gap closing, suggests a topological phase transition associated with increasing interlayer coupling strengths.
To determine the topological phases, we calculate the quadrupole moment $q_{xy}$ [25] using the nested Wilson loop. As a result, we find that $q_{xy}=\frac{1}{2}$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{bilayer}}$, identifying the system as topological, while $q_{xy}=0$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{bilayer}}$, indicating a trivial phase. These results indicate that the interlayer coupling induces a topological phase transition in the BBH bilayer system. The presence and absence of topological corner states changes at the critical value $t_{\mathrm{c}}^{\rm{bilayer}}$ as the band gap closes and re-opens. This critical value marks the transition point where the number of topological corner states changes. The detailed calculation of the quadrupole moment is shown in Appendix C.
II.2 Topological corner states in BBH trilayer system type $\mathrm{I}$ (topological-trivial-topological structure)
We proceed to study a BBH trilayer system depicted in Fig. 1(e). This arrangement consists of a trivial layer positioned between two topological layers, where the layers are coupled by interlayer coupling $t_{\mathrm{v}}$.
@@
with the interlayer coupling matrix $\mathcal{H}_{\text{v}}$. Similarly to the bilayer system, the labeling of sites in the Hamiltonian varies between odd and even layers to maintain the chiral symmetry, as shown in Fig. 1(d). This labeling is crucial for applying the methods we will discuss in the next section. The trilayer system has six pairs of doubly degenerate bands. Similarly to the bilayer system, the band gap closes at a critical value $t_{\mathrm{c}}^{\rm{trilayer}}=\kappa_{1}+\kappa_{2}$ (see Appendix A). The band gap reopens again when $t_{\mathrm{v}}$ exceeds the critical value.
@@
Figure 4: (a) Energy spectrum vs $t_{\mathrm{v}}$ for BBH trilayer system type $\mathrm{I}$ (topological-trivial-topological structure). Blue dots represent the gapped edge states and bulk spectrum. The green and red lines correspond to two and a single set of corner modes, respectively. (b) Field amplitude summed over the four corner modes at $E=0$ for $t_{\mathrm{v}}$=$t_{\mathrm{c}}^{\rm{trilayer}}-1$ and $t_{\mathrm{v}}$=$t_{\mathrm{c}}^{\rm{trilayer}}+1$. ’Corner-asym’ persists even after exceeding the critical value, while ’Corner-sym’ vanishes at the critical value $t_{\mathrm{c}}^{\rm{trilayer}}$. The inset shows an enlarged view of a single corner, with the gray area representing a unit cell. Parameters are chosen as $\kappa_{1}$=1, $\kappa_{2}=3$, and $N=20$.
@@
For a finite BBH trilayer system with square geometry, with each layer consisting of $2N\times 2N$ sites, we calculate the eigenenergies as a function of the interlayer coupling strength(Fig 4(a)). When $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$, which is below the critical value, there are two sets of corner states, with a total of eight (illustrated by the green line in the figure). In contrast, when $t_{\mathrm{v}}$ exceeds $t_{\mathrm{c}}^{\rm{trilayer}}$, only a single set of corner states survives, with a total number of four (indicated by the red line). Each of these sets consists of four degenerate modes at $E=0$. These two sets of corner states are distinguished by their unique field distributions(Fig. 4 (b)). Specifically, the corner states that disappear at the critical interlayer coupling show non-zero field amplitudes across all layers, which corresponds to symmetric coupling between the corners (Corner-sym). Conversely, the corner states that remain after exceeding the critical interlayer coupling exhibit a complete absence of field amplitude in the middle layer, which corresponds to anti-symmetric coupling between the corners (Corner-asym). The detailed analysis of the field distribution for the corner states under the $\kappa_{1}=0$ limit is provided in Appendix B. At large interlayer couplings, the presence of topological corner states, ’Corner-asym’, is explained through an effective model near zero energy, derived based on the perturbation theory. See Appendix D.
@@
As discussed in subsection $\mathrm{II.A}$ of the bilayer system, we calculate the quadrupole moment $q_{xy}$ to identify the topological phases of the trilayer system. Consequently, our analysis shows that $q_{xy}=0$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$, classifying the system as trivial, despite the presence of two sets of corner states at $E=0$. In the case where $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$, we have $q_{xy}=\frac{1}{2}$, indicating the system as topological. This leads to the issue where the quadrupole moment fails to fully capture the topological phases in certain instances. Therefore, the introduction of another invariant is essential to accurately describe the topological phases of both BBH bilayer and trilayer systems. We will calculate and discuss this alternative invariant in the next section.
II.3 Topological corner states in BBH trilayer system type $\mathrm{II}$ (trivial-topological-trivial structure)
@@
Next, we study a trilayer system that involves a topological layer sandwiched between two trivial layers, each layer coupled by an interlayer coupling $t_{\mathrm{v}}$. This arrangement is depicted in Fig. 1 (f). In this case, the Hamiltonian is given by
@@
Figure 5: (a) Energy spectrum vs $t_{\mathrm{v}}$ for BBH trilayer system type $\mathrm{II}$ (trivial-topological-trivial structure). Blue dots correspond to the bulk spectrum, with or without gapped edges, whereas the red line represents a single set of corner modes. Gapped edges are present when the system exhibits corner states. (b) Field amplitude summed over the four corner modes at $E=0$ for $t_{\mathrm{v}}$=$t_{\mathrm{c}}^{\rm{trilayer}}-1$. The inset shows an enlarged view of a single corner, with the gray area representing a unit cell. Parameters are chosen as $\kappa_{1}$=3, $\kappa_{2}=1$ and $N=20$.
Fig. 5(a) shows the energy spectrum as a function of $t_{\mathrm{v}}$. There is a single set of corner modes that includes four degenerate corner states at $E=0$ (depicted by the red line) when $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$; however, all corner modes disappear above the critical value $t_{\mathrm{c}}^{\rm{trilayer}}$. An enlarged view of a single mode reveals that the corner mode distribution is localized on the sublattice opposite to that of the topological-trivial-topological structure. See Appendix B for details. The absence of topological corner states at large interlayer couplings can be explained by an effective model near zero energy, derived through perturbation theory, as detailed in Appendix D.
The field amplitudes of the corner modes, shown in Fig. 5 (b), are distributed across each layer. The quadrupole moments are obtained as $q_{xy}=\frac{1}{2}$ for $t_{\mathrm{v}}<t_{\mathrm{c}}^{\rm{trilayer}}$, identifying the phase as topological, while $q_{xy}=0$ for $t_{\mathrm{v}}>t_{\mathrm{c}}^{\rm{trilayer}}$, classifying it as trivial. We note that in BBH trilayer systems with large interlayer coupling $t_{\mathrm{v}}$, the size of the band-gap varies between two distinct systems, despite the fact that they have identical band structures, as shown in Figs. 4(a) and 5(a). This discrepancy is caused by the presence of gapped edge states within the band gap, as illustrated in Fig. 4(a).